# CoBRA: A Universal Strategyproof Confirmation Protocol for Quorum-based Proof-of-Stake Blockchains

### v2

@@ 4.1 CoBRA Overview
  • Liveness. A reward scheme that incentivizes the participation of rational validators and mitigates censorship attacks.
@@
  • Efficiency. A strong chain mechanism that enables higher-volume finalization during periods of high participation.
4.2 SMR protocol
We model the SMR protocol $\Pi$ as a black-box abstraction that proceeds in rounds. In some rounds, one or more validators may propose blocks. Validators can vote on proposed blocks either explicitly, by casting votes, or implicitly, e.g., by extending a block with a new proposal. A block is considered confirmed in $\Pi$ once a validator obtains a valid quorum certificate for it. Whenever a validator $p$ obtains a valid certificate $QC(b)$ for block $b$ at height $h$, the protocol outputs to $p$: deliver$(QC(b),b,h)$.
This abstraction captures a broad class of $q$-committable blockchain protocols, including leader-based, chained, DAG-based, and leaderless designs. For instance, in leader-based protocols, designated leaders propose blocks in certain rounds, followed by one or more voting rounds leading to block finalization.
We classify the action space of validators during the execution of $\Pi$ into two categories: safety-preserving and liveness-preserving actions.
Safety-Preserving Actions
Rational validators may collude with byzantine validators to fork the system. More specifically, although the SMR protocol $\Pi$ may output a single block $b$ with a valid certificate $QC(b)$ to $p_{i}$, before the message delay elapses, rational and Byzantine validators can produce a conflicting block $b^{\prime}$ with its own valid certificate $QC(b^{\prime})$ by obtaining signatures from correct validators that have not yet witnessed $QC(b)$.
We assume a broadcast-based SMR protocol $\Pi$ [^3] Otherwise, validators in our construction can broadcast the valid certificates to propagate the required information., ensuring that all conflicting blocks with valid certificates are detected by every correct validator within $2\Delta^{*}$  [^4] Clients only accept blocks with $12f+$ 1 finality votes according to the client confirmation rule. Validators commit finality votes only for blocks with valid certificates that meet conditions specified by the protocol. Hence, rational validators cannot benefit from hiding votes from correct validators. Therefore, we consider only certificates witnessed by at least one correct validator.. We further analyze this attack in Appendix A.
Liveness-Preserving Actions
Since $\Pi$ is $(n,f)$-resilient, it guarantees liveness as long as there are at most $f$ Byzantine validators and at least $n-f$ correct validators. Given that the set of correct and rational validators is $f+1+k\geq n-f$, the protocol remains live if the $k$ rational validators participate in liveness-preserving actions. In the setting of $q$–committable blockchains, however, a valid certificate is sufficient to commit a block. Therefore, it is enough to guarantee that there exist specific intervals during which rational validators participate in the liveness-preserving steps.
Liveness-preserving actions include (i) proposing a valid block when it is their turn according to $\Pi$, (ii) voting for valid blocks proposed by others, and (iii) engaging in additional actions such as view-change protocols specified by $\Pi$. Assuming synchronous periods are sufficiently long, these conditions guarantee that at least one correct validator will eventually propose and finalize a valid block. As a result, any valid transaction observed by all correct validators will eventually be included in the ledger.
In contrast, rational validators may choose to abstain from liveness-preserving actions if offered a sufficiently high bribe by the adversary to censor specific transactions. We further elaborate on liveness attacks in Appendix A and provide examples of bribing mechanisms in Section 7.
4.3 Finality Gadget
The finality gadget ensures that validators finalize only transactions that cannot be reversed and enables them to provide inclusion proofs to clients.
@@ 4.3 Finality Gadget
We present the finality gadget in Algorithm 1. Moreover, below, we describe the validator and client protocols for accepting a transaction. Since clients do not monitor the protocol execution, their confirmation rule differs from that of validators. As a result, validators may finalize transactions based on protocol guarantees, whereas clients rely on additional proofs to verify finality.
@@
When a validator $p_{i}$ witnesses a valid certificate $QC(b)$ for a block $b$, it stores the certificate (line 7) along with the block and the current local timestamp (line 8). Then $p$ initiates the block finalization function.
Since conflicting blocks may exist but remain unwitnessed within the last $2\Delta^{*}$, validators employ the Stake-bounded finalization rule. More specifically, each validator $p$ traverses its local ledger and using the timestamp of the blocks and its local clock, determines which blocks are $2\Delta^{*}$–old (output $2\Delta^{*}$ ago), or $2\Delta^{*}$–recent (output within the last $\Delta^{*}$).  [^5] To measure elapsed time, $p_{i}$ relies on its local clock rather than requiring synchronized clocks (lines 8, 18). Then, $p$ finalizes blocks as follows.
$2\Delta^{*}$–old Blocks
If no conflicting block has been detected, validators finalize any block $b^{\prime}$ that was output by $\Pi$ at least $2\Delta^{*}$ ago (lines 19 and 20). If a conflicting block existed, it would have been observed by this time.
$2\Delta^{*}$–recent Blocks
For blocks output within the last $2\Delta^{*}$, validators cannot guarantee the absence of conflicting chains. Instead, each validator finalizes only a subset of the most recent blocks, ensuring that the total finalized transaction volume does not exceed $C=D$ coins (lines 22 and 23). This guarantees that any rational validator engaging in an attack can gain at most $D$ coins. For intuition, we provide a tight example in Figure 1, where the adversary partitions all correct validators to create $f+1$ conflicting ledgers before the bound $\Delta^{*}$ elapses. Assuming Byzantine validators claim no rewards and instead use them to bribe the $f$ rational validators they need to successfully double-spend, the $f$ rational validators share a total of $f\times D$ double-spent coins, resulting in a per-validator gain of $D$.
Figure 1: Forking attack.
At round $t$, a subset $\mathcal{F}_{R}$ of $2f$ misbehaving validators (Byzantine and rational) fork the system. The adversary partitions correct validators into $f+1$ distinct sets and collects their signatures to create conflicting blocks with valid certificates. Each correct validator finalizes blocks with transaction volume up to $C=D$ within the last $\Delta^{*}$. The misbehaving rational validators share the rewards and, consequently, each rational validator is able to double-spend $D$ coins.
@@
After finalizing a block $b$, validators include finality votes for $b$ in subsequent blocks. A block $b$ is considered finalized for clients once at least $12f+$ 1 finality votes for $b$ or a block extending $b$ have been recorded.
To confirm the inclusion of a transaction $tx$, validators provide clients with a proof consisting of:
@@
  • A finality proof containing at least $12f+$ 1 finality votes for $b$ or a descendant block.
This ensures that clients can securely verify the transaction’s finalization, ensuring that the SMR protocol is certifiable.
4.4 Rewards and Penalties
@@
To ensure misbehaving validators are penalized, we implement a two-stage withdrawal process [16]:
@@
  • Withdrawals are only allowed after a waiting period $\Delta_{W}>2\Delta^{*}$, during which validators cannot participate in the protocol.
@@
  • If a validator is detected signing conflicting blocks, it is blacklisted (line 13), preventing permanently its withdrawal.
All correct validators will detect the equivocations within $2\Delta^{*}$. Thus, misbehaving validators cannot reclaim their stake. This disincentivizes rational validators from forking.
Recovery
If a fork is detected, honest validators halt participation in the underlying SMR protocol $\Pi$. As a result, the system stalls due to an insufficient number of active validators to ensure liveness. However, as we will prove, forking is not a profitable strategy for rational validators and thus does not occur in equilibrium. We briefly discuss recovery and reconfiguration procedures following misbehavior, such as rational validators misestimating their utility, in Appendix 7.
Algorithm 1 Finality Gadget
@@ 8       $ledger\leftarrow ledger\cup(b,localtime)$
11       $(block,\_\;)\leftarrow ledger[h]$
12       if $block\neq b$ then
13          slash($quorumstore[h]\cap QC(b)$)
14          $quorumstore[h].add(QC(b))$
15
16procedure finalize-blocks
@@ 17    $suffixvalue=0$
19       if $blocktime\leq localtime-2\Delta^{*}\land|blacklist|=\emptyset$ then .
@@ 21          $suffixindex++$
22       else if $suffixvalue+block.value\leq C\land|blacklist|=\emptyset$ then

### v3

@@ 4. CoBRA: Finalization Rule
  • A reward scheme that incentivizes the participation of rational validators and mitigates censorship attacks (liveness). Transaction fees are distributed to participating validators only in execution windows in which at least nfn-f distinct validators successfully contribute to block production.
@@
  • A strong chain mechanism that enables higher-volume finalization during periods of high participation, improving efficiency.
4.1. Finality Gadget
@@
The finality gadget ensures that validators finalize only transactions that cannot be reversed and enables them to provide inclusion proofs to clients. Below, we describe the validator and client protocols for accepting transactions. Since clients do not monitor the protocol execution, their confirmation rule differs from that of validators. As a result, validators may finalize transactions based on protocol guarantees, whereas clients rely on additional proofs to verify finality.
@@
We assume a $q$-committable SMR protocol $\Pi$ that is broadcast-based [^4] Otherwise, validators in our construction can broadcast the valid certificates to propagate the required information. to ensure that safety violations are detected by every correct validator within $\Delta^{*}$. We treat $\Pi$ as a black-box abstraction: whenever a validator obtains a valid quorum certificate $QC(b)$ for a block $b$ at height $h$, the protocol outputs $\mathsf{deliver}(QC(b),b,h)$ to that validator (line 10). Then, the validator stores the certificate (line 12) along with the block and the current local timestamp (line 13), and initiates the block finalization function.
Since conflicting blocks may exist but remain unwitnessed within the last $\Delta^{*}$, validators employ the stake-bounded finalization rule. More specifically, each validator $p$ traverses its local ledger and using the timestamp of the blocks and its local clock, determines which blocks are $\Delta^{*}$–old (output $\Delta^{*}$ ago or more), or $\Delta^{*}$–recent (output within the last $\Delta^{*}$) [^5] To measure elapsed time, the validator relies on its local clock rather than requiring synchronized clocks (lines 13, 36).. Then, $p$ finalizes blocks as follows.
$\Delta^{*}$–old Blocks
If no conflicting block has been detected, validators finalize any block $b^{\prime}$ that was output by $\Pi$ at least $\Delta^{*}$ ago (lines 37 and 38).
$\Delta^{*}$–recent Blocks
For blocks output within the last $\Delta^{*}$, each validator finalizes only a subset of the most recent blocks, ensuring that the total finalized transaction volume does not exceed $C=D$ coins (lines 41 and 42). This guarantees that any rational validator engaging in an attack can gain at most $D$ coins. For intuition, we provide a tight example in Figure 2 for $f^{*}=f$ Byzantine and $k=f$ rational validators.
Figure 2. Forking attack.
A subset $\mathcal{F}_{R}$ of $2f$ misbehaving validators (Byzantine and rational) fork the system. The adversary partitions all correct validators into $f+1$ distinct sets and collects their signatures to create conflicting blocks with valid certificates before $\Delta^{*}$ elapses. Each correct validator finalizes blocks with transaction volume up to $C=D$ within $\Delta^{*}$. Assuming Byzantine validators claim no rewards and instead use them to bribe the $f$ rational validators they need to successfully double-spend, then the $f$ rational validators share a total of $f\times D$ double-spent coins, resulting in a per-validator gain of $D$.
Finality votes
To enable inclusion proofs and ensure accountability, validators include finality votes for a block $b$ in subsequent blocks once $b$ is finalized. We say that $b$ has a finality certificate, denoted $\mathcal{F}(b)$, if at least $2f+1$ finality votes for $b$ (or for any block extending $b$) are recorded on-chain, with each vote included in a block carrying a valid quorum certificate.
Conflicting certificates.
Correct validators never commit finality votes for conflicting blocks. If a validator observes conflicting blocks with valid certificates or finality certificates (lines 17, 29), it immediately stops finalizing new blocks (lines 37, 41). By limiting the transaction volume so that the total potential gain for any rational validator does not exceed their stake, we can ensure that, once they are slashed, rational validators have no incentive to engage in equivocations.
@@
Clients, who do not participate directly in the consensus protocol, cannot observe when validators’ internal confirmation rules are met. The finality certificate thus serves as a verifiable proof of finality, allowing clients to confirm that a block has been finalized according to the validators’ rules.
Specifically, to confirm the inclusion of a transaction $tx$, validators provide clients with a proof consisting of:
@@
  • A finality certificate $\mathcal{F}(b)$ for the block $b$, including the finality votes and the valid quorum certificates of the blocks including these votes.
In this way, clients can securely verify transaction finalization, ensuring that the SMR protocol is certifiable.
4.2. Rewards and Penalties
@@
To ensure misbehaving validators are penalized, we implement a two-phase withdrawal process (Org, 2024):
@@
  • Withdrawals are only allowed after a period $\Delta_{W}>2\Delta^{*}$, during which validators cannot participate in the protocol.
@@
  • If a validator is detected signing finality votes for conflicting blocks, it is blacklisted (line 30), preventing permanently its withdrawal.
Note that the withdrawal period is set to be larger than $2\Delta^{*}$ (instead of $\Delta^{*}$). Although equivocations are observed within time $\Delta^{*}$—which is sufficient to limit the transaction volume—only after $2\Delta^{*}$ can we be certain that no conflicting block exists. For example, consider a block finalized at time tt [^6] We can fix the local clock of any validator, as only the elapsed time matters. by $2f$ rational and Byzantine validators along with one honest validator $p_{i}$. Just before time $t+\Delta^{*}$, the $2f$ malicious validators may finalize a conflicting block with the help of another honest validator $p_{j}$. In the worst case, $p_{i}$ will become aware of the conflict only at time $t+2\Delta^{*}$. Setting the extended withdrawal period larger than $2\Delta^{*}$ is sufficient to prevent malicious validators from withdrawing their funds prematurely.
Rewards
Participation rewards are sourced from transaction fees and distributed to validators based on their contributions to block finalization. Rewards are proportional to the fraction of blocks with valid certificates proposed by each validator and are distibuted every $n$ rounds. Specifically, we define the interval of rounds $[kn+1,(k+1)n]$ for $k\in\mathbb{Z}$ as the leader window $W_{k}$. After the end of the window $W_{k}$ the total transaction fees from finalized transactions $F_{W_{k}}$ are allocated based on the proportion of unique proposers $P_{W_{k}}$ relative to $n$: $R_{W_{k}}=F_{{W_{k}}}\cdot\dfrac{P_{W_{k}}}{n}.$ To ensure that correct validators successfully propose blocks, rewards are distributed only if $P_{W_{k}}\geq n-f$, i.e., at least nfn-f unique validators submit proposals within $n$ rounds.
This guarantees censorship resilience, as Byzantine validators can only earn rewards in windows that include at least one honest proposer. Even if a Byzantine validator attempts to participate normally to accumulate rewards in order to bribe rational validators later, they would eventually deplete their coins. With no coins left to bribe, rational validators prefer to simply follow the protocol and collect the participation rewards.
4.3. Extending Finalization with Strong Chains
Recall that validators finalize transaction volumes up to $D$ coins within a time window $\Delta^{*}$. This is the lower bound in our threat model. When the system is not under attack (i.e., fewer faulty nodes), blocks often receive more than $2f+1$ signatures. We leverage the certificates with additional signatures to further limit forks and finalize larger transaction volumes without waiting for $\Delta^{*}$.
To achieve this, we extend the concept of $i$-strong chains (Xiang et al., 2021) in our model. A ledger is $i$-strong if it contains valid certificates signed by $2f+1+i$ unique validators within $\Delta^{*}$. With at least $1+i$ correct validators, the maximum number of forks decreases to fif-i, allowing secure finalization of higher transaction volumes. In particular, when a certificate has $2.5f+1$ signatures, we can define a strongest chain that can finalize infinite transaction volume (regardless of $D$).
Definition 0 ($i$-strong chain).
Consider a validator $p$ and its local data structure $\text{ledger}^{p}$ (Algorithm 1, line 1). Let $S$ represent the set of distinct signatures from validators included in the valid certificates of blocks within the set of blocks in $\text{ledger}^{p}$ that were output by the SMR protocol $\Pi$ within the last $\Delta^{*}$. We say that $\text{ledger}^{p}$ is an $i$-strong chain, where i=|S|2f1i=|S|-2f-1.
Strong chain finalization rule
Strongest chain finalization rule
We additionally define the strongest chain as an i-strong chain with $i>\frac{f+1}{2}$. More than half of the correct validators participate in the strongest chain. This ensures that only one strongest chain can exist within $\Delta^{*}$, enabling correct validators to finalize every transaction, under this rule,
5. CoBRA: Recovery Mechanism
In Section 4, we describe CoBRA’s finalization rule which implements an $(n,k,f^{*})$–resilient transformation for executions with $f^{*}\leq f$ Byzantine validators and $k\leq 2f-f^{*}$ rational validators. In this regime, although equivocations are possible with the participation of rational parties, they are not profitable for rational validators; thus, rational parties will not engage in forking attacks.
We now extend our protocol to executions with $f+1\leq f^{*}<5n/9$ Byzantine validators, and $k<2n/3-f^{*}$ rational validators. In this setting, equivocations cannot be prevented, since $f+1$ Byzantine validators alone suffice to create blocks with conflicting certificates within $\Delta^{*}$. Accordingly, we no longer aim to preclude safety violations; instead, we provide a recoverable economic-security regime that guarantees economic restitution with recovery parameter $\Delta_{R}=4\Delta^{*}$ (Definition 2.6).
Overview.
At a high level, CoBRA’s recovery protocol proceeds in the following steps:
  1. (1)
    Detecting conflicts: Upon observing conflicting certificates, correct validators halt further finalization to prevent additional financial damage to clients. Validators learn all client-acceptable finalized blocks by disseminating the relevant certificates containing finality votes.
  2. (2)
    Execution and reimbursement: Validators revert to a stable pre-fork state and deterministically re-execute all conflicting finalized ledgers in a common order, reimbursing provable client losses from the stake of provably misbehaving validators, without minting new coins.
  3. (3)
    New genesis: Validators issue a verifiable commitment to the recovered state, allowing the system to resume from a single consistent ledger.
5.1. Detecting Conflicts
We first bound the maximum economic damage that clients may incur and then describe how validators reliably observe all conflicting finalized blocks following an equivocation.
Economic damage
Clients accept only blocks with finality certificates. According to the finalization rules of CoBRA, the total amount of coins carrying finality votes during a fork is at most $fD$. Once the synchronous bound elapses and correct validators are aware of the equivocation they stop finalizing new blocks.
Observing conflicting certificates.
Ensuring that honest validators will witness all finalized blocks to reimburse affected clients is not straightforward. Recall that each vote in a finality certificate is included in a block with a valid SMR quorum certificate (namely an ss–phase certificate [^7] By the definition of $q$-commitable SMR protocols, a quorum certificate is formed by collecting $q$ votes across $s$ phases (e.g., $s=2$ in PBFT, $s=3$ in HotStuff).) which is broadcast to all validators. Nevetheless, malicious validators may withhold their votes for the ss–phase certificate from other validators while sending these votes directly to clients. As a result, a block may be accepted by clients even though honest validators have not observed its finality certificate.
To ensure that every block accepted by clients is eventually executed by validators, without requiring clients to disseminate certificates, validators will re-execute any block that carries $2f+1$ finality votes, provided that each vote is included in a block with at least an (s1)(s-1)–phase certificate. Whenever a block carries a valid ss–phase certificate whose finality votes form a finality certificate, at least one honest validator must have witnessed the corresponding (s1)(s\!-\!1)–phase certificates and finality votes.
Thus, upon detecting the first such violation, validators broadcast all (s1)(s\!-\!1)–phase certificates containing finality votes that form valid finality certificates but for which no corresponding ss–phase certificate has been observed (line 52). They then wait an additional $2\Delta^{*}$ (line 53) to ensure that all conflicting finalized ledgers have been revealed, as detailed in Section 4.2 (slashing mechanism). Thus, within $2\Delta^{*}$, all validators are guaranteed to have received the complete set $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$, where $j\leq f+1$, which consists of all conflicting blocks whose $2f+1$ finality votes appear in at least (s1)(s\!-\!1)–phase certificates [^8] The withholding attack described above can happen only for the last block of any conflicting ledger in $\mathcal{L}$. For every intermediate block, the existence of an extending block implies that its ss–phase certificate was witnessed by an honest validator. Nevertheless, validators proactively broadcast all (s1)(s\!-\!1)–phase certificates upon equivocation to avoid waiting an additional $2\Delta^{*}$ until knowing the last block per ledger..
5.2. Execution and Reimbursement
We now describe how validators execute transactions, including those in the conflicting ledgers $\mathcal{L}$, to update the system state. Recovering a common state requires that all validators execute the same transactions in the same order, even in the presence of conflicting finalized ledgers. To this end, validators revert to the state up to a block for which no conflicting finalized block exists. To support this efficiently, validators locally maintain two distinct states.
  • Recent-finalized state: the state up to their most recent finalized block,
  • Old-finalized state: the state up to their most recent $2\Delta^{*}$–old finalized block (i.e., output $2\Delta^{*}$ ago).
Upon detecting an equivocation, validators revert to their old-finalized state (line 56), for which it is guaranteed that no fork exists. After waiting $2\Delta^{*}$, during which all conflicting finalized blocks in $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$ are observed, validators update the system state.
To do so, validators first re-execute all transactions from their old-finalized state up to the first point of conflict, namely the first block in $\mathcal{L}=\{L_{1},\ldots,L_{j}\}$ (lines 56-59). For the rest blocks, they sort the conflicting ledgers in $\mathcal{L}$ lexicographically by the hash of their last finalized block, forming a new ordered set $\mathcal{L}^{*}$ (line 55). Finally, validators execute each ledger in $\mathcal{L}^{*}$ sequentially, processing all block within each ledger in order (lines 60-62).
During execution, transactions fall into two categories. First, transactions issued by honest clients who believed a branch to be canonical remain valid, since honest clients do not spend more coins that they hold. Second, transactions from malicious participants attempting to double-spend. For these, validators reimburse affected clients from the misbehaving validators’ slashed stake and burn any remaining stake to punish the malicious validators (line 63).
Algorithm 1 CoBRA protocol
@@ 13   $ledger\leftarrow ledger\cup(b,localtime)$
16   $(block,\_\_\;)\leftarrow ledger[h]$
17   if $block\neq b$ then $violation\leftarrow true$
18
19on observing (“finality”, $u,b$) in a phase–(s1)(s-1) certificate do
20  finality_votes[b].add(u)finality\_votes[b].add(u)
21  if $|finality\_votes[b]|\geq 2f+1\land finalitystore[b.height]=\bot$ then
22   $finalized.apply(b)$ $\triangleright$ Execute $b$ to update the state
23   $(\_\_,time)\leftarrow ledger[b.height]$
24   if $localtime-time\geq 2\Delta^{*}\land b.height>oldFinalizedBlock.height$ then
25    $oldFinalized.apply(b)$, $oldFinalizedBlock\leftarrow b$
26  else if $|finality\_votes[b]|\geq 2f+1\land finalitystore[b.height]\neq\bot$ then
27   $(b^{\prime},\_\_)\leftarrow finalitystore[b.height]$
28   if $b^{\prime}\neq b$ then
29    $violation\leftarrow true$
30    $blacklist\leftarrow blacklist\cup(finalitystore[b.height]\cap finality\_votes[b])$
31    if this is the first time conflicting blocks are observed, Recover
32    finalitystore[b.height].add(finality_votes[b])finalitystore[b.height].add(finality\_votes[b])
33
34procedure finalize-blocks
@@ 35  $suffixvalue=0$
37   if $blocktime\leq localtime-\Delta^{*}\land\;(not\;violation)$ then
@@
39    broadcast signed finality vote $v$ for block: ("finality", $v$, block)
@@ 40    $suffixindex++$
41   else if $suffixvalue+block.value\leq C\land\;(not\;violation)$ then
@@
43    broadcast signed finality vote $v$ for block: ("finality", $v$, block)